Circular No. 6175

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

SUPERNOVA 1995P IN ANONYMOUS GALAXY R. H. McNaught, Anglo-Australian Observatory, reports his discovery

of a supernova (mag ∼ 18) found on a J plate taken by M. J. Drinkwater with the U.K. Schmidt Telescope (UKST) on May 6.5 UT and located at α = 14^{h}07^{m}14^{s}.33, δ = −2^{o}43′32′′.7 (equinox 2000.0; uncertainty 0′′.3 in each coordinate), which is 1′′.3 east and 4′′.6 north of the galaxy’s center. No star appears to mag ∼ 19 on a 3-min R film taken by Q. A. Parker with the UKST on 1995 Feb. 27.7. A nearby star of mag 17 has position end figures 14^{s}.17, 07′′.5.

B. Schmidt, Mount Stromlo and Siding Spring Observatories (MSSSO), communicates that a spectrogram was taken with the MSSSO 2.3-m telescope, revealing broad features that are most consistent with a type-Ia supernova ∼ 6 weeks past maximum; other types can not be unequivocally ruled out due to the very poor signal of the spectrum.

S. Benetti and M. Turatto, European Southern Observatory, write: “We have observed SN 1995P on May 30.2 UT with the New Technology Telescope (+ EMMI) at La Silla under non-photometric conditions. A spectrum (range 380–846 nm, resolution 1 nm) closely resembles those of the type-Ia SN 1994D, ∼ 1 month past maximum light (cf. Patat et al. MNRAS, in press). In particular, the slopes of the continua are similar, indicating that both supernovae suffered the same amount of reddening. Broad emissions are measured at 488, 511, 523, 585, 616, and 688 nm. SN 1995P is located 1′′.1 east and 4′′.7 north of the nucleus of a lenticular parent galaxy. Its redshift, determined from Ca II (H and K), Mg I, and Na I absorption lines, is 0.056.”

NOVA AQUILAE 1995 Further to their astrometry on IAUC 6174, R. W. Argyle and L. V.

Morrison, Royal Greenwich Observatory, report photometry obtained with the Carlsberg Automatic Meridian Circle. Following are measured V magnitudes (±0.08 mag): May 9.22 UT, 11.59; 13.21, 11.74; 15.20, 11.75; 18.19, 11.90; 20.19, 11.84; 21.18, 11.85.

Visual magnitude estimates: May 21.958, 11.6 (T. Vanmunster, Landen, Belgium); 24.032, 11.8 (P. Schmeer, Bischmisheim, Germany); 26.018, 11.8 (L. Szentaskó, Veresegyház, Hungary).

1995 May 30 Daniel W. E. Green